# Totally geodesic submanifolds of the homogeneous nearly Kähler 6-manifolds and their $\mathsf{G}_{2}$-cones

### v2

@@ 5. The examples
Many of the totally geodesic submanifolds that will appear in this section are isometric to a sphere with a round or complex Berger metric. We can compute the radius $r$ of the sphere $\mathsf{S}^{n}(r)$, as well as the parameters of the Berger sphere $\mathsf{S}^{3}_{\mathbb{C},\tau}(r)$ from its sectional curvature. Indeed, it is well known that the sectional curvature of $\mathsf{S}^{n}(r)$ is equal to $1/r^{2}$. In the case of $\mathsf{S}^{3}_{\mathbb{C},\tau}(r)$, the parameters $r$ and $\tau$ can be obtained from the equations $\tau=r^{2}\sec(U,X)$ and $4-3\tau=r^{2}\sec(X,Y)$, where $U$ is a vertical vector and $X$, $Y$ are horizontal vectors with respect to the Hopf fibration (see [21]).
@@ 5.1. The complex projective space
We describe the totally geodesic examples of the complex projective space $\mathbb{C}\mathsf{P}^{3}$ equipped with a homogeneous nearly Kähler metric.
@@
$\mathfrak{u}(2)^{j}=\operatorname{span}\{jE_{11},jE_{22},E_{21}-E_{12},j(E_{12}+E_{21})\}=(\mathfrak{u}(2)^{j}\cap\mathfrak{k})\oplus(\mathfrak{u}(2)^{j}\cap\mathfrak{p}).$
@@ 5.1.3. The horizontal sphere $\mathsf{SU}(2)\cdot o$
$\mathfrak{su}(2)=\operatorname{span}\{i(E_{11}-E_{22}),E_{21}-E_{12},i(E_{12}+E_{21})\}=(\mathfrak{su}(2)\cap\mathfrak{k})\oplus(\mathfrak{su}(2)\cap\mathfrak{p}),$
@@ 5.1.4. The sphere $\mathsf{SU}(2)_{\Lambda_{3}}\cdot o$
$\mathfrak{su}(2)_{\Lambda_{3}}=\operatorname{span}\left\{i(E_{11}+3E_{22}),\sqrt{2}e_{1}+\sqrt{3}e_{3},\sqrt{2}e_{2}+\sqrt{3}e_{4}\right\}=(\mathfrak{su}(2)_{\Lambda_{3}}\cap\mathfrak{k})\oplus(\mathfrak{su}(2)_{\Lambda_{3}}\cap\mathfrak{p}),$
@@ 5.2. The flag manifold
We describe the totally geodesic examples of the flag manifold $\mathsf{F}(\mathbb{C}^{3})$ equipped with a homogeneous nearly Kähler metric.
5.2.1. The real flag manifold $\mathsf{F}(\mathbb{R}^{3})$ [54, Example 3.1]
There is a natural embedding of the real flag manifold $\mathsf{F}(\mathbb{R}^{3})$ in $\mathsf{F}(\mathbb{C}^{3})$ which is induced by the usual inclusion of $\mathbb{R}^{3}$ in $\mathbb{C}^{3}$. This submanifold can also be seen as the orbit $\mathsf{SO}(3)\cdot o$ of the standard $\mathsf{SO}(3)\subseteq\mathsf{SU}(3)$, and the corresponding isotropy subgroup is $\mathsf{SO}(3)_{o}=\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$, so we get $\mathsf{F}(\mathbb{R}^{3})=\mathsf{SO}(3)/\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}=\mathsf{Sp}(1)/\mathsf{Q}_{8}$, where $\mathsf{Q}_{8}=\{\pm 1,\pm i,\pm j,\pm k\}$. Observe that $\mathfrak{so}(3)$ is canonically embedded in $\mathfrak{su}(3)$, since $\mathfrak{so}(3)\subseteq\mathfrak{p}$. Thus, Theorem 4.3 allows us to conclude that $\mathsf{F}(\mathbb{R}^{3})$ is totally geodesic in $M$, and its tangent space is precisely $\mathfrak{p}_{\mathsf{F}(\mathbb{R}^{3})}=\mathfrak{so}(3)=\operatorname{span}\{e_{1},e_{3},e_{5}\}$. A direct computation shows that $\mathsf{F}(\mathbb{R}^{3})$ has constant curvature equal to $1/8$. Furthermore, we have the equality $J(\mathfrak{so}(3))=\operatorname{span}\{e_{2},e_{4},e_{6}\}$, implying that the inclusion $\mathsf{F}(\mathbb{R}^{3})\subseteq\mathsf{F}(\mathbb{C}^{3})$ is Lagrangian. Finally, note that $\mathsf{F}(\mathbb{R}^{3})$ is well-positioned at $o$, so $\mathsf{F}(\mathbb{R}^{3})$ is well-positioned by Corollary 4.10.
5.2.2. The Berger sphere [54, Example 3.2]
@@
Consider the maximal torus $\mathsf{H}\subseteq\mathsf{SU}(3)$ whose Lie algebra is given by $\mathfrak{h}=\operatorname{span}\{e_{1}+e_{3}+e_{5},e_{2}+e_{4}-e_{6}\}$. Observe that $\mathfrak{h}\subseteq\mathfrak{p}$, so it is a canonically embedded subalgebra of $\mathfrak{g}$, and the orbit $\mathsf{H}\cdot o$ is a totally geodesic surface. Since $\mathsf{H}\cdot o$ is a quotient of $\mathsf{H}$ by a finite group, it is a compact abelian Lie group itself and hence diffeomorphic to a torus. However, we will see that $\mathsf{H}\cdot o$ is not isometric to the standard flat torus. Indeed, in order to determine the isometry type of $\mathsf{H}\cdot o$, we compute the preimage $\exp_{o}^{-1}(o)$. For this, we will need a description of the Riemannian exponential map $\exp_{o}\colon\mathfrak{h}\to\mathsf{H}\cdot o\subseteq M$, which is merely the restriction of the Riemannian exponential map of $M$. To this end, we define the orthonormal vectors $X=\frac{1}{\sqrt{3}}(e_{1}+e_{3}+e_{5})$, $Y=\frac{1}{\sqrt{3}}(e_{2}+e_{4}-e_{6})$ of $\mathfrak{h}$. Then the exponential map of $\mathfrak{h}$ satisfies that $\exp_{o}(uX+vY)=e^{uX}e^{vY}\cdot o$, and this element is equal to $o$ if and only if $e^{uX}e^{vY}$ is a diagonal matrix. The solutions to $\exp_{o}(uX+vY)=o$ are given by the lattice $\Lambda=\operatorname{span}_{\mathbb{Z}}\left\{\left(\sqrt{2}\pi,\sqrt{2}\pi/\sqrt{3}\right),\left(0,2\sqrt{2}\pi/\sqrt{3}\right)\right\}$. Since $\exp_{o}$ is $\mathfrak{h}$-equivariant, in the sense that it satisfies the equation $\exp_{o}(T+S)=\operatorname{Exp}(T)\cdot\exp_{o}(S),$ it follows that $\exp_{o}$ is actually a Riemannian covering map, so $\mathsf{H}\cdot o$ is isometric to the quotient $\mathbb{R}^{2}/\Lambda$. We will refer to this orbit as $\mathsf{T}^{2}_{\Lambda}=\mathsf{H}\cdot o=\mathbb{R}^{2}/\Lambda$. Note that $\mathsf{H}\cdot o$ is not a product $\mathsf{S}^{1}(r_{1})\times\mathsf{S}^{1}(r_{2})$, since the closest points in $\Lambda\setminus\{(0,0)\}$ to the origin are those in {(±$\left\{\left(\pm\sqrt{2}\pi,\pm\sqrt{6}\pi/3\right),\left(\pm\sqrt{2}\pi,\mp\sqrt{6}\pi/3\right),\left(0,\pm 2\sqrt{6}\pi/3\right)\right\},$ and thus there exist three different closed geodesics of minimum length, as opposed to two in the case of $\mathsf{S}^{1}(r)\times\mathsf{S}^{1}(r)$ or one in the case of $\mathsf{S}^{1}(r_{1})\times\mathsf{S}^{1}(r_{2})$ with $r_{1}\neq r_{2}$. Let $\mathfrak{p}_{\mathsf{T}^{2}_{\Lambda}}=\mathfrak{h}=\operatorname{span}\{e_{1}+e_{3}+e_{5},e_{2}+e_{4}-e_{6}\}$ be the tangent space of this surface, then $J(\mathfrak{p}_{\mathsf{T}^{2}_{\Lambda}})=\mathfrak{p}_{\mathsf{T}^{2}_{\Lambda}}$, and by homogeneity it follows that $\mathsf{T}^{2}_{\Lambda}$ is an almost complex surface in $\mathsf{F}(\mathbb{C}^{3})$. However, it is clear from the expression of $\mathfrak{p}_{\mathsf{T}^{2}_{\Lambda}}$ that $\mathsf{T}^{2}_{\Lambda}$ is not well-positioned.
5.2.4. The fiber of the submersion $\mathsf{F}(\mathbb{C}^{3})\to\mathbb{C}\mathsf{P}^{2}$ [13, Example 3.1]
@@
is a subgroup of $\mathsf{SU}(3)$ conjugate to the standard $\mathsf{SO}(3)$. The corresponding Lie algebra is given by $\mathfrak{so}(3)^{\sigma}=\operatorname{span}\{\operatorname{diag}(i,0,-i),e_{1}+e_{3},e_{2}+e_{4}\}$, and in particular it is canonically embedded in $\mathfrak{su}(3)$. One sees that the isotropy subgroup $\mathsf{SO}(3)^{\sigma}_{o}$ is the $\mathsf{U}(1)$ subgroup generated by $\mathfrak{so}(3)^{\sigma}\cap\mathfrak{t}$, so we obtain that $\mathsf{SO}(3)^{\sigma}\cdot o$ is a totally geodesic submanifold of $\mathsf{F}(\mathbb{C}^{3})$ that is diffeomorphic to a sphere. Its tangent space at $o$ is $\mathfrak{p}_{\mathfrak{so}(3)^{\sigma}}=\operatorname{span}\{e_{1}+e_{3},e_{2}+e_{4}\},$ and this plane has sectional curvature $1/2$, so $\mathsf{SO}(3)^{\sigma}\cdot o$ is isometric to a two-dimensional sphere of radius $\sqrt{2}$. The equality $J(\mathfrak{p}_{\mathfrak{so}(3)^{\sigma}})=\mathfrak{p}_{\mathfrak{so}(3)^{\sigma}}$ implies that $\mathsf{SO}(3)^{\sigma}\cdot o$ is an almost complex surface in $\mathsf{F}(\mathbb{C})^{3}$. Since $\mathsf{SO}(3)^{\sigma}\cdot o$ is not well-positioned at $o$, it is not well-positioned.
@@ 5.2.6. Real projective planes inside $\mathsf{F}(\mathbb{R}^{3})$
We view $\mathbb{H}\equiv\mathbb{R}^{4}$. Consider the totally geodesic embedding $h\colon\mathsf{S}^{2}(2\sqrt{2})\hookrightarrow\mathsf{S}^{3}(2\sqrt{2})$ defined by $h(x,y,z)=xi+yj+zk$. Then $\pi\circ h$ is also a totally geodesic immersion of the $3$-sphere satisfying $\pi\circ h(-x,-y,-z)=\pi\circ h(x,y,z)$ for all $(x,y,z)\in\mathsf{S}^{2}(2\sqrt{2})$, so it factors through an isometric immersion $\phi\colon\mathbb{R}\mathsf{P}^{2}(2\sqrt{2})\to\mathsf{F}(\mathbb{R}^{3})$ defined via
@@ 5.3. The almost product $\mathsf{S}^{3}\times\mathsf{S}^{3}$
Let $\Sigma=\mathsf{S}^{3}$ be the fiber of the projection map $(x,y)\mapsto x$, which we know from Subsection 2.2.3 that it is a totally geodesic submanifold of $\mathsf{S}^{3}\times\mathsf{S}^{3}$, and it is coincides with the orbit $\left(\Delta_{1,3}\mathsf{SU}(2)\times\mathsf{SU}(2)_{2}\right)\cdot o$. It is immediate to check that the normalizer of $\Sigma$ in $\mathsf{G}$ is precisely $\mathsf{N}_{\mathsf{G}}(\Sigma)=\Delta_{1,3}\mathsf{SU}(2)\times\mathsf{SU}(2)_{2}$, and the restricted action $\mathsf{N}_{\mathsf{G}}(\Sigma)\curvearrowright\Sigma$ satisfies
@@ 5.3.3. The torus [8, Example 1]
so the corresponding lattice is $\Gamma=\operatorname{span}_{\mathbb{Z}}\left\{(2\pi/\sqrt{3},2\pi),(4\pi/\sqrt{3},0)\right\}$. Thus, $\mathsf{T}^{2}\cdot o$ is isometric to the flat torus $\mathsf{T}^{2}_{\Gamma}=\mathbb{R}^{2}/\Gamma$. It turns out that the closest points in $\Gamma\setminus\{(0,0)\}$ to the origin are those in the set {±$\left\{\pm\left(2\pi/\sqrt{3},2\pi\right),\pm\left(4\pi/\sqrt{3},0\right),\pm\left(2\pi/\sqrt{3},-2\pi\right)\right\}$, so $\mathsf{T}^{2}_{\Gamma}$ admits three closed geodesics of minimum length $\frac{4\pi}{\sqrt{3}}$, unlike any Riemannian product of the form $\mathsf{S}^{1}(a)\times\mathsf{S}^{1}(b)$. By construction, we have that $\mathfrak{p}_{\mathsf{T}^{2}_{\Gamma}}=\operatorname{span}\{e_{1},e_{4}\}$ is a $J$-invariant subspace, so $\mathsf{T}^{2}_{\Gamma}$ is an almost complex surface inside $\mathsf{S}^{3}\times\mathsf{S}^{3}$. By Corollary 4.10, $\mathsf{T}^{2}_{\Gamma}$ is a well-positioned totally geodesic submanifold of $\mathsf{S}^{3}\times\mathsf{S}^{3}$.

### v3

@@ 5. The examples
Many of the totally geodesic submanifolds that appear in this section are isometric to a sphere with a round or complex Berger metric. We can compute the radius $r$ of the sphere $\mathsf{S}^{n}(r)$, as well as the parameters of the Berger sphere $\mathsf{S}^{3}_{\mathbb{C},\tau}(r)$ from its sectional curvature. Indeed, it is well known that the sectional curvature of $\mathsf{S}^{n}(r)$ is equal to $1/r^{2}$. In the case of $\mathsf{S}^{3}_{\mathbb{C},\tau}(r)$, the parameters $r$ and $\tau$ can be obtained from the equations $\tau=r^{2}\sec(U,X)$ and $4-3\tau=r^{2}\sec(X,Y)$, where $U$ is a vertical vector and $X$, $Y$ are horizontal vectors with respect to the Hopf fibration (see [21]).
@@ 5.1. The complex projective space
We describe the totally geodesic examples of the nearly Kähler complex projective space $\mathbb{C}\mathsf{P}^{3}$.
@@
$\mathfrak{u}(2)^{j}=\operatorname{span}\bigl{\{}jE_{11},jE_{22},E_{21}-E_{12},j(E_{12}+E_{21})\bigr{\}}=(\mathfrak{u}(2)^{j}\cap\mathfrak{k})\oplus(\mathfrak{u}(2)^{j}\cap\mathfrak{p}).$
@@ 5.1.3. The horizontal sphere $\mathsf{SU}(2)\cdot o$
$\mathfrak{su}(2)=\operatorname{span}\bigl{\{}i(E_{11}-E_{22}),E_{21}-E_{12},i(E_{12}+E_{21})\bigr{\}}=(\mathfrak{su}(2)\cap\mathfrak{k})\oplus(\mathfrak{su}(2)\cap\mathfrak{p}),$
@@ 5.1.4. The sphere $\mathsf{SU}(2)_{\Lambda_{3}}\cdot o$
$\mathfrak{su}(2)_{\Lambda_{3}}=\operatorname{span}\Bigl{\{}i(E_{11}+3E_{22}),\sqrt{2}e_{1}+\sqrt{3}e_{3},\sqrt{2}e_{2}+\sqrt{3}e_{4}\Bigr{\}}=(\mathfrak{su}(2)_{\Lambda_{3}}\cap\mathfrak{k})\oplus(\mathfrak{su}(2)_{\Lambda_{3}}\cap\mathfrak{p}),$
@@ 5.2. The flag manifold
We describe the totally geodesic examples of the nearly Kähler flag manifold $\mathsf{F}(\mathbb{C}^{3})$.
5.2.1. The real flag manifold $\mathsf{F}(\mathbb{R}^{3})$ [53, Example 3.1]
There is a natural embedding of the real flag manifold $\mathsf{F}(\mathbb{R}^{3})$ in $\mathsf{F}(\mathbb{C}^{3})$ which is induced by the usual inclusion of $\mathbb{R}^{3}$ in $\mathbb{C}^{3}$. This submanifold can also be seen as the orbit $\mathsf{SO}(3)\cdot o$ of the standard $\mathsf{SO}(3)\subseteq\mathsf{SU}(3)$, and the corresponding isotropy subgroup is $\mathsf{SO}(3)_{o}=\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$, so we get $\mathsf{F}(\mathbb{R}^{3})=\mathsf{SO}(3)/(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})=\mathsf{Sp}(1)/\mathsf{Q}_{8}$, where $\mathsf{Q}_{8}=\{\pm 1,\pm i,\pm j,\pm k\}$. Observe that $\mathfrak{so}(3)$ is canonically embedded in $\mathfrak{su}(3)$, since $\mathfrak{so}(3)\subseteq\mathfrak{p}$. Thus, Theorem 4.3 allows us to conclude that $\mathsf{F}(\mathbb{R}^{3})$ is totally geodesic in $M$, and its tangent space is precisely $\mathfrak{p}_{\mathsf{F}(\mathbb{R}^{3})}=\mathfrak{so}(3)=\operatorname{span}\{e_{1},e_{3},e_{5}\}$. A direct computation shows that $\mathsf{F}(\mathbb{R}^{3})$ has constant curvature equal to $1/8$. Furthermore, we have the equality $J(\mathfrak{so}(3))=\operatorname{span}\{e_{2},e_{4},e_{6}\}$, implying that the inclusion $\mathsf{F}(\mathbb{R}^{3})\subseteq\mathsf{F}(\mathbb{C}^{3})$ is Lagrangian. Finally, note that $\mathsf{F}(\mathbb{R}^{3})$ is well-positioned at $o$, so $\mathsf{F}(\mathbb{R}^{3})$ is well-positioned by Corollary 4.10.
5.2.2. The Berger sphere [53, Example 3.2]
@@
Consider the maximal torus $\mathsf{H}\subseteq\mathsf{SU}(3)$ whose Lie algebra is given by $\mathfrak{h}=\operatorname{span}\{e_{1}+e_{3}+e_{5},e_{2}+e_{4}-e_{6}\}$. Observe that $\mathfrak{h}\subseteq\mathfrak{p}$, so it is a canonically embedded subalgebra of $\mathfrak{g}$, and the orbit $\mathsf{H}\cdot o$ is a totally geodesic surface. Since $\mathsf{H}\cdot o$ is a quotient of $\mathsf{H}$ by a finite group, it is a compact abelian Lie group itself and hence diffeomorphic to a torus. However, we see that $\mathsf{H}\cdot o$ is not isometric to the standard flat torus. Indeed, in order to determine the isometry type of $\mathsf{H}\cdot o$, we compute the preimage $\exp_{o}^{-1}(o)$. For this, we need a description of the Riemannian exponential map $\exp_{o}\colon\mathfrak{h}\to\mathsf{H}\cdot o\subseteq M$, which is merely the restriction of the Riemannian exponential map of $M$. To this end, we define the following orthonormal vectors $X=\frac{1}{\sqrt{3}}(e_{1}+e_{3}+e_{5})$, $Y=\frac{1}{\sqrt{3}}(e_{2}+e_{4}-e_{6})$ of $\mathfrak{h}$. Then the exponential map of $\mathfrak{h}$ satisfies $\exp_{o}(uX+vY)=e^{uX}e^{vY}\cdot o$, and this element is equal to $o$ if and only if $e^{uX}e^{vY}$ is a diagonal matrix. The solutions to $\exp_{o}(uX+vY)=o$ are given by the lattice $\Lambda=\operatorname{span}_{\mathbb{Z}}\left\{\left(\sqrt{2}\pi,\sqrt{2}\pi/\sqrt{3}\right),\left(0,2\sqrt{2}\pi/\sqrt{3}\right)\right\}$. Since $\exp_{o}$ is $\mathfrak{h}$-equivariant, in the sense that it satisfies the equation $\exp_{o}(T+S)=\operatorname{Exp}(T)\cdot\exp_{o}(S),$ it follows that $\exp_{o}$ is actually a Riemannian covering map, so $\mathsf{H}\cdot o$ is isometric to the quotient $\mathbb{R}^{2}/\Lambda$. We will refer to this orbit as $\mathsf{T}^{2}_{\Lambda}=\mathsf{H}\cdot o=\mathbb{R}^{2}/\Lambda$. Note that $\mathsf{H}\cdot o$ is not a product $\mathsf{S}^{1}(r_{1})\times\mathsf{S}^{1}(r_{2})$, since the closest points in $\Lambda\setminus\{(0,0)\}$ to the origin are those in {(±$\left\{\left(\pm\sqrt{2}\pi,\pm\sqrt{6}\pi/3\right),\left(\pm\sqrt{2}\pi,\mp\sqrt{6}\pi/3\right),\left(0,\pm 2\sqrt{6}\pi/3\right)\right\},$ and thus there exist three different closed geodesics of minimum length passing through $o$, as opposed to two in the case of $\mathsf{S}^{1}(r)\times\mathsf{S}^{1}(r)$ or one in the case of $\mathsf{S}^{1}(r_{1})\times\mathsf{S}^{1}(r_{2})$ with $r_{1}\neq r_{2}$. Let $\mathfrak{p}_{\mathsf{T}^{2}_{\Lambda}}=\mathfrak{h}=\operatorname{span}\{e_{1}+e_{3}+e_{5},e_{2}+e_{4}-e_{6}\}$ be the tangent space of this surface, then $J(\mathfrak{p}_{\mathsf{T}^{2}_{\Lambda}})=\mathfrak{p}_{\mathsf{T}^{2}_{\Lambda}}$, and by homogeneity it follows that $\mathsf{T}^{2}_{\Lambda}$ is an almost complex surface in $\mathsf{F}(\mathbb{C}^{3})$. However, it is clear from the expression of $\mathfrak{p}_{\mathsf{T}^{2}_{\Lambda}}$ that $\mathsf{T}^{2}_{\Lambda}$ is not well-positioned.
5.2.4. The fiber of the submersion $\mathsf{F}(\mathbb{C}^{3})\to\mathbb{C}\mathsf{P}^{2}$ [15, Example 3.1]
@@
is a subgroup of $\mathsf{SU}(3)$ conjugate to the standard $\mathsf{SO}(3)$. The corresponding Lie algebra is given by $\mathfrak{so}(3)^{\sigma}=\operatorname{span}\{\operatorname{diag}(i,0,-i),e_{1}+e_{3},e_{2}+e_{4}\}$, and in particular it is canonically embedded in $\mathfrak{su}(3)$. One sees that the isotropy subgroup $\mathsf{SO}(3)^{\sigma}_{o}$ is the $\mathsf{U}(1)$ subgroup generated by $\mathfrak{so}(3)^{\sigma}\cap\mathfrak{k}$, so we obtain that $\mathsf{SO}(3)^{\sigma}\cdot o$ is a totally geodesic submanifold of $\mathsf{F}(\mathbb{C}^{3})$ that is diffeomorphic to a sphere. Its tangent space at $o$ is $\mathfrak{p}_{\mathfrak{so}(3)^{\sigma}}=\operatorname{span}\{e_{1}+e_{3},e_{2}+e_{4}\},$ and this plane has sectional curvature $1/2$, so $\mathsf{SO}(3)^{\sigma}\cdot o$ is isometric to a two-dimensional sphere of radius $\sqrt{2}$. The equality $J(\mathfrak{p}_{\mathfrak{so}(3)^{\sigma}})=\mathfrak{p}_{\mathfrak{so}(3)^{\sigma}}$ implies that $\mathsf{SO}(3)^{\sigma}\cdot o$ is an almost complex surface in $\mathsf{F}(\mathbb{C})^{3}$. Since $\mathsf{SO}(3)^{\sigma}\cdot o$ is not well-positioned at $o$, it is not well-positioned.
@@ 5.2.6. Real projective planes inside $\mathsf{F}(\mathbb{R}^{3})$
We view $\mathbb{H}\equiv\mathbb{R}^{4}$. Consider the totally geodesic embedding $h\colon\mathsf{S}^{2}(2\sqrt{2})\hookrightarrow\mathsf{S}^{3}(2\sqrt{2})$ defined by $h(x,y,z)=xi+yj+zk$. Then $\pi\circ h$ is also a totally geodesic immersion of the $2$-sphere satisfying $\pi\circ h(-x,-y,-z)=\pi\circ h(x,y,z)$ for all $(x,y,z)\in\mathsf{S}^{2}(2\sqrt{2})$, so it factors through an isometric immersion $\phi\colon\mathbb{R}\mathsf{P}^{2}(2\sqrt{2})\to\mathsf{F}(\mathbb{R}^{3})$ defined via
@@ 5.3. The almost product $\mathsf{S}^{3}\times\mathsf{S}^{3}$
Let $\Sigma=\mathsf{S}^{3}$ be the fiber of the projection map $(x,y)\mapsto x$, which we know from Subsection 2.2.3 that is a totally geodesic submanifold of $\mathsf{S}^{3}\times\mathsf{S}^{3}$, and coincides with the orbit $\left(\Delta_{1,3}\mathsf{SU}(2)\times\mathsf{SU}(2)_{2}\right)\cdot o$. It is immediate to check that the normalizer of $\Sigma$ in $\mathsf{G}$ is precisely $\mathsf{N}_{\mathsf{G}}(\Sigma)=\Delta_{1,3}\mathsf{SU}(2)\times\mathsf{SU}(2)_{2}$, and the restricted action $\mathsf{N}_{\mathsf{G}}(\Sigma)\curvearrowright\Sigma$ satisfies
@@ 5.3.3. The torus [8, Example 1]
so the corresponding lattice is $\Gamma=\operatorname{span}_{\mathbb{Z}}\left\{(2\pi/\sqrt{3},2\pi),(4\pi/\sqrt{3},0)\right\}$. Thus, $\mathsf{T}^{2}\cdot o$ is isometric to the flat torus $\mathsf{T}^{2}_{\Gamma}=\mathbb{R}^{2}/\Gamma$. It turns out that the closest points in $\Gamma\setminus\{(0,0)\}$ to the origin are those in the set {±$\left\{\pm\left(2\pi/\sqrt{3},2\pi\right),\pm\left(4\pi/\sqrt{3},0\right),\pm\left(2\pi/\sqrt{3},-2\pi\right)\right\}$, so $\mathsf{T}^{2}_{\Gamma}$ admits three closed geodesics of minimum length $\frac{4\pi}{\sqrt{3}}$ passing through $o$, unlike any Riemannian product of the form $\mathsf{S}^{1}(a)\times\mathsf{S}^{1}(b)$. By construction, we have that $\mathfrak{p}_{\mathsf{T}^{2}_{\Gamma}}=\operatorname{span}\{e_{1},e_{4}\}$ is a $J$-invariant subspace, so $\mathsf{T}^{2}_{\Gamma}$ is an almost complex surface inside $\mathsf{S}^{3}\times\mathsf{S}^{3}$. By Corollary 4.10, $\mathsf{T}^{2}_{\Gamma}$ is a well-positioned totally geodesic submanifold of $\mathsf{S}^{3}\times\mathsf{S}^{3}$.